Ahmet is 26 years young er than his mother. In 4 years, his mother will be 3 times Ahmet's age. How old is Ahmet now?
Answer by JulietG(1812)

M = A + 26
M + 4 = 3(A+4)
M + 4 = 3A + 12
M = 3A + 8
from the top equation:
A + 26 = 3A + 8
A + 18 = 3A
18 = 2A
9 = A
If Ahmet is 9, then his mother is 35.
In 4 years, Ahmet will be 13 and his mother will be 39.